OpenStudy (anonymous):

Please help. Find the surface area of the sphere which has a diameter of 30. A. 188.5 B. 706.86 C. 2827.43 D. 14137.17 I am not a math person AT ALL.

OpenStudy (badhi):

surface area of a sphere is, $$4\pi r^2$$ where r is the radius knowing that diameter (d) =2r you can easily find the surface area http://en.wikipedia.org/wiki/Sphere

OpenStudy (anonymous):

4πr2 x = 4π15^2 x = 12.55637(225) x = 2827.433388

OpenStudy (badhi):

yes, you're correct

OpenStudy (anonymous):

Thank you both so much. I know it might be easy to find, but I am really horrible at math. Thanks!!!!!!

OpenStudy (anonymous):

haha anytime ;)